Is (-2, 3) a solution of the system? Explain why or why not in a complete sentence.

y = 2x + 7
x = y - 4

Answer by jim_thompson5910(35256)

No it is not because it does not satisfy the equation x = y-4.

x = y-4
-2 = 3-4 ... plug in (x,y) = (-2,3)
-2 = -1 ... this is false

To be a solution to the system, it needs to satisfy BOTH equations. It needs to make both equations true. If one is false, then we don't even need to bother check the other equation.